OpenStudy (anonymous):

Calculus help please…? Im lost on this one :/ find f'(x) and f'(c) f(x)=x+1/x-1 c=2

OpenStudy (anonymous):

quotient rule for this one \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x+1,f'(x)=1,g(x)=x-1,g'(x)=1\]

OpenStudy (anonymous):

=(1)(x-1)+(1)(x+1)???

OpenStudy (anonymous):

@satellite73